OpenStudy (goldendragon01):

will medal 8{58-37(58-6)}-3

OpenStudy (goldendragon01):

can u help

OpenStudy (anonymous):

first do (58-6)=52 than 37(52)=1924 than 58-1924=-1866 than 8(-1866)= - 14928 than -14928-3=14931

OpenStudy (anonymous):

its -14931